A 0.79-kg ball is placed on top of a 2.5-kg book which is at the top of a table. What is the magnitude of the force exerted by the table to the two objects?
1
Expert's answer
2020-12-14T07:32:57-0500

Let's apply the Newton's Second Law of Motion:


"\sum F_y=ma_y=0,""N_{table}-m_{book}g-m_{ball}g=0,""N_{table}=(m_{book}+m_{ball})g,""N_{table}=(2.5\ kg+0.79\ kg)\cdot 9.8\ \dfrac{m}{s^2}=32.24\ N."

Answer:

"N_{table}=32.24\ N."